# Learn the Ropes, Then Trust the Wins: Self-imitation with Progressive Exploration for Agentic Reinforcement Learning

### v2

@@ 4.2 Intrinsic Reward Shaping
To provide a strong baseline for the agentic performance after RL, we refer to the existing studies [85, 86, 87] for the enhancement of diverse exploration, stable convergence, and effective training. Our baseline, termed as Dr.BoT, consists of the following bag-of-tricks modifications to the vanilla GRPO [39].
Removal of KL Divergence
Clip-Higher
Removal of Intra-group Normalization
Removal of Length Normalization
Filtering of Over-long and Void-turn Samples
Filtering of Low-variance Groups
5 Experiments
In this section, we present the empirical evaluations of our SPEAR across a variety of LLM agent tasks.
5.1 Experimental Setup
Environments
In accordance with our design of agent action space (see Section 3.2), we train the LLM agents on three challenging benchmarks: ALFWorld [5], WebShop [42], and DAPO-Math-17K [46]. ALFWorld is an interactive environment created to evaluate how well LLM agents can handle multi-step decision-making tasks. In each scenario, the agent is given a textual goal and must achieve it by engaging in multiple rounds of interaction with the environment. The platform offers 4,639 task examples spanning six typical household activity categories: Pick & Place (Pick), Examine in Light (Look), Clean & Place (Clean), Heat & Place (Heat), Cool & Place (Cool), and Pick Two & Place (Pick2). WebShop, on the other hand, is a sophisticated web-based platform aimed at assessing LLM agents in authentic online shopping situations. Agents are required to interact with a simulated HTML shopping site to search for products, browse items, and purchase an appropriate product. WebShop supports a broad and varied action space, featuring more than 1.1 million products and 12K user instructions. DAPO-Math-17K is a rigorously engineered, competition-grade benchmark designed to stress-test large-scale RL on LLM agents. The agent must develop multi-step mathematical reasoning, perform strategic tool-calling for code verification, and reflect on feedback from the sandbox before submitting its final answer. It contains 17K manually-curated prompts sourced from olympiad-level problems, each transformed so that every ground-truth label is an integer—eliminating symbolic-parsing noise and yielding a clean, deterministic reward signal.
Baselines
With respect to ALFWorld and WebShop, we follow [41] to report a range of competitive baselines such as prompting-based method (i.e., direct I/O) for the proprietary models GPT-4o [88] and Gemini [89], framework-based method such ReAct [4] and Reflexion [90], RL methods including PPO [20], RLOO [91, 92], GRPO [39, 2], GiGPO [41], and our proposed strong baseline Dr.BoT. For the DAPO-Math-17K, we follow [34] to report the performance of domain-specific expert (e.g., Qwen2.5-Math [93]), existing reasoning models (e.g., Sky-T1 [94], o1 [45], DeepSeek-distilled Qwen 32B [2], QwQ [95], and s1 [96]), and the tool-integrated RL counterparts (e.g., ReTool [34], SimpleTIR [18], ZeroTIR [31], and AFM [97]).
Training Details
For ALFWorld and WebShop, we follow [41] to use Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct [93] as our base models. For DAPO-MATH-17K, we follow [34] to use Qwen2.5-32B-Instruct [93] for fair comparison. In addition, we use the latest Qwen3-32B-Instruct [98] for generalization studies. All the training settings and the details on hyper-parameters can be found in the appendix Section A.2. We use VeRL [99] as our codebase for RL.
5.2 Performance
Table 1: Performance on ALFWorld and WebShop (%). For ALFWorld, we report the average success rate for each subtask as well as the overall results. For WebShop, we report the average score and the success rate (SR). PT, FW, and RL stand for prompting, framework, and reinforcement learning, respectively.
@@ 5.2 Performance
Table 2: Performance (mean@30) on AIME 2024 and AIME 2025 (%). : Results of the official ReTool implementation under the VeRL framework are cited where the GRPO already utilizes multiple DAPO tricks for faster convergence and better performance. : Official results reported by Qwen [98]. PT and RL stand for prompting and reinforcement learning, respectively.
@@
PT I/O Qwen2.5-Math-72B 4K 30.0
PT I/O Qwen2.5-Math-72B CI 4K 40.0
PT I/O Sky-T1 32K 43.3
PT I/O o1-preview 128K 44.6 37.9
PT I/O DeepSeek-R1-Qwen-32B 32K 47.0
PT I/O QwQ-32B-Preview 32K 50.0 33.5
PT I/O QwQ-32B 38K 79.5 69.5
PT I/O s1-32B 32K 56.7
@@
Table 1 demonstrates the effectiveness of our proposed SPEAR on both ALFWorld and WebShop. We show that the SPEAR is compatible with various GRPO-like baselines including the vanilla GRPO [39], the competitive GiGPO [41], and our proposed strong baseline Dr.BoT. SPEAR brings consistent performance gains over baselines across 1.5B and 7B models up to 20%. We believe such generalization benefits from the collection of successful trajectories during training, which acts as a walkthrough guide to the agent. Especially for these two tasks where the success rate is fairly low at the beginning, the agent has to figure out the underlying logic of interaction with the environment (e.g., the simulated world and the offline website), and summarizes from the trajectories for its own action paradigm tailored specific to the tasks. In this case, the replayed successful trajectories expedite the accumulation of promising reasoning-action experiences, and thereafter reduce blind trials and errors under the given budget of interaction turns. Furthermore, our proposed Dr.BoT boosts the performance of either GRPO and GiGPO up to 15%, showcasing that the mixture of tricks indeed benefits the trade-off between exploration and exploitation, and stabilizes agentic training on long-horizon tasks.
Table 2 reports the performance under the settings of CI-integrated reasoning for solving math problems. As shown, the proposed Dr.BoT indeed outperforms a series of recent RL methods as a strong baseline. It is noted that the strict comparability among existing RL methods cannot be guaranteed due to the following reasons: 1) the dataset used for training, 2) the implementation framework, and 3) the hyper-parameter settings. Especially for the context length, we observe that the maximum number of tokens allowed during inference exists a great effect on the reported results. For example, the results of pure-text reasoning on Qwen3 are reported under the context of 38K and the reduced context length impedes answer parsing from the overlong responses. Given the CI tool, the agentic LLM learns to exploit the feedback from the sandbox for double-check and self-reflection. It achieves comparable performance on Qwen3 but with a much smaller token budget, highlighting the efficiency of reasoning with code. When the context is relaxed to 32K during inference, an improvement is observed on both Qwen2.5 and Qwen3, confirming the generalization of our SPEAR with more interactions turns and reasoning tokens. Due to the long reasoning nature of Qwen3, its performance with and without our SPEAR on the AIME24 remain almost the same under 16K budget where over 10% trajectories are overlong. With 32K tokens, our superiority is revealed on both AIME24 and AIME25.
@@ 5.3 Ablation Study
Table 3: Ablation study on ALFWorld and WebShop. For ALFWorld, we report the average success rate (%) for each subtask as well as the overall result. For WebShop, we report both the average score and the average success rate (SR) (%). SI and IR stand for Self-Imitation and Intrinsic Reward, respectively.
@@
Table 4: Ablation study (mean@30) on AIME 2024 and AIME 2025 (%). : Results of the official ReTool implementation under the VeRL framework are cited where the GRPO already utilizes multiple DAPO tricks for faster convergence and better performance. SI and IR stand for Self-Imitation and Intrinsic Reward, respectively.
@@
To investigate the effectiveness of our SPEAR, we conduct ablation studies on various tasks and baselines. Tables 3 and 4 respectively report the results on ALFWorld and WebShop, and DAPO-MATH-17K.
Self-Imitation
@@
The SIL improves over baselines consistently on the ALFWorld and WebShop across model scales. Since either 1.5B or 7B models perform poorly at the early stage (i.e., success rate $15\%$), past experiences are quite beneficial to LLM agents for exploring promising strategies. The re-use of the precious trajectories facilitates convergence and prevents mechanical trials especially for small LLMs. For DAPO-MATH-17K, the performance on AIME24 dropped a bit by self-imitation but that on AIME25 still gets improved. Such fluctuation is related to the phenomenon we observed in Figure 5 where the imitation of samples with multiple tool calls leads to rapid increase of interaction turns and thereafter causes training instability. The competition between different reward terms affects the robust selection of correct, good experiences, ultimately degrading the effectiveness of self-imitation.
Intrinsic Reward
@@
For the ALFWorld and WebShop tasks, we observe that the rewarding of interaction turns benefit 1.5B models consistently. For 7B models, two outliers are found where the self-imitation alone brings the most performance gains (up to 13%). Such exception might be related to both the task definition and the RL algorithm in use. One should experiment with different combinations in practice to determine the optimal settings. For DAPO-MATH-17K, we find that the intrinsic reward is indispensable for both Qwen2.5 and Qwen3 models because the reward to encourage more interactions with the sandbox is critical to the transform from text-based reasoning to tool-integrated reasoning. Both the two models have been pretrained on massive mathematic corpus and tend to solve maths problems with textual reasoning. The intrinsic reward promotes frequent tool calling and such rich observation signals from the sandbox motivate the agent to correct coding errors, check the validity of the answer, and reflect on alternative solutions.
(a) Replay Buffer Size $N_{\mathcal{D}}$.
(b) Baseline Buffer Size $N_{\mathcal{D}_{R}}$.
(c) Covariance-based Clipping Ratio $\lambda$.
(d) Warm-up Steps $T_{\text{warm-up}}$.
(e) Decay Steps $T_{\text{decay}}$.
5.4 Discussions on Hyper-parameters
Replay Buffer Size $N_{\mathcal{D}}$
Baseline Buffer Size $N_{\mathcal{D}_{R}}$
When $N_{\mathcal{D}_{R}}=0$, the original advantages are used without recalibration and filtering (see Equation 8). It shows that the direct imitation of these experiences can be suboptimal where certain trajectories are outdated for the current policy. By timely adjusting the advantages and removing inappropriate experiences ($\tilde{A}_{j}\leq 0$), we reduce the inaccurate estimation for off-policy update. We also find that $N_{\mathcal{D}_{R}}$ should not be set too large as such 50-th percentile reward deviates from the latest ones, decreasing the effectiveness of recalibration.
Covariance-based Clipping Ratio $\lambda$
Warm-up Step $T_{\text{warm-up}}$
Decay Step $T_{\text{decay}}$
5.5 Qualitative Analysis
(a) Before RL training.
(b) After RL training.
(c) The evolution of efficient coding from the purpose of computation to verification.
Figure 8: Development of the agent’s coding skills.
Skill Development
We follow [34] to analyze the coding capabilities of the agent before and after RL by classifying the purpose of the code snippets. Specifically, we employ Hunyuan-Large [100] to interpret reasoning contexts before each tool-calling and judge the intention of the codes passed into the code interpreter on DAPO-MATH-17K dataset. The code purposes with their frequency over twice are kept and illustrated in Figure 8. We find that the after RL, the agent becomes proficient in writing codes for solution finding, problem-solving, and solution verification, which is quite advanced compared with the basic calculation and computation before RL. Our case study shows that after training, the agent learns to master matrix computation with numpy for higher efficiency, confirming the improvement of coding skills.
Figure 9: The advancement of strategy from the search query perfectionism to goal-oriented active progression.
Action Refinement
5.6 Training Cost and Complexity
Training
Stage
Computation
of GRPO
Additional
Computation
by SPEAR
Description
RB Update $\mathcal{O}(GT)$ Copy operation (negligible).
On-policy BP $\mathcal{O}(P)$ BP w/ $\pi_{\theta}$.
Replay PG
$n_{\text{SIL}}KT\mathcal{O}(M)$
+ $n_{\text{SIL}}\mathcal{O}(KT)$
FW w/ $\pi_{\theta}$, token-wise
$\text{clip}\&\min$ (negligible).
Replay BP $n_{\text{SIL}}\mathcal{O}(P)$ BP w/ $\pi_{\theta}$.
(a) ALFWorld.
(b) WebShop.
(c) DAPO-MATH-17K.
@@ 5.7 Generalization on Vision-Language Agents
Figure 11: During training, the agentic LLM learns to push the box towards the target for reward.
To test whether the proposed SPEAR is still complimentary to existing GRPO-like algorithms on training visual agents, we follow [41] to conduct experiments on the popular visual game Sokoban [102]. In this setting, the Qwen2.5-VL-3B-Instruct [103] is adopted as the agentic LLM to solve the puzzle game where the player must push the boxes along the grid towards target positions without hitting the walls. It challenges the agent on spatial comprehension and long-term planning capabilities. The grid size is of $6\times 6$ and the visual agent receives both the visual (RGB arrays) and textual inputs as states. As shown in Table 6, the proposed method generally improves the performance on Sokoban with either GRPO, GiGPO, and the proposed Dr.BoT baselines. At first, the visual agent is unaware of the winning logic behind the game and wanders around for ”aimlessly” exploration (see Figure 11(a)). After optimization, it not only comprehends the spatial relationship to control the box but also learns to stop moving when the task is completed.
@@ 6 Conclusions and Limitations
In this paper, we target the pivotal challenge of balancing exploration and exploitation in RL training of LLM agents. Our proposed solution, SPEAR [Uncaptioned image], extends the vanilla SIL by incorporating a replay buffer of high-return trajectories with advantage recalibration, scheduled entropy control, and intrinsic rewards. These components work in a curriculum manner to prevent policy collapse and excessive uncertainty, progressively guiding the policy through a smooth transition between exploration and exploitation. In addition, we propose a strong baseline Dr.BoT tailored for agentic RL training with existing bag-of-tricks verified from numerical industrial practices. Empirical results across various tasks and models showcase SPEAR’s superiority over existing methods, with performance boosts and acceptable computational overhead. The effectiveness of our SPEAR underscores the value of learning from past experiences while managing policy entropy, offering a robust framework for training LLMs with strong reasoning and tool integration skills.
There exist two potential limitations:
  • The vague definition of good experiences under highly complex, stochastic environments with unreliable tools. In such cases, the observations from the tool (e.g., online searching) or the environments (e.g., simulation of robotics, traffics, and financial trading) can be noisy and severely degrade the feasibility of the task. The sparse outcome reward cannot distinguish between good and bad experiences and therefore the relative advantages might be simply attributed to randomness instead of the agent’s behavior (e.g., reasoning capacity or action rationality). We suggest a possible solution that more fine-grained, stepwise supervision should be enforced. For example, for each tool response or the environment feedback, an LLM-based judge determines whether the contents are meaningful for the final solution. In addition, a step-wise process reward that evaluates the logical consistency [84] of the agent’s thought and action might be helpful in selecting good experiences (although it may also lead to negative outcomes).
  • The rigidity of entropy control which relies on prior-based scheduling and covariance-based clipping. In the present study, we design regularization techniques based on our prior knowledge about policy entropy as a proxy to monitor exploration and exploitation situations. Such scheduling and clipping designs might not be optimal for all kinds of agentic tasks. We believe a more adaptive solution lies in the policy’s self-confidence on decisions under the observations of each task. For example, one might use the token-level dynamic reweighting for the SIL loss [104] which avoids over-concentration on certain low-probability reference tokens in the replay buffer. Similarly, the covariance-based clipping can also introduce the token-wise output probability into the sampling process to replace the vanilla bounded random sampling.
We leave the exploration mentioned above as a promising direction for improvement in the future.
@@ Appendix A Appendix
A.1 Pseudo Code
Algorithm 1 summarizes the full training procedure of the proposed SPEAR. It is noted that our SPEAR is compatible with various baselines such as GRPO [39] and GiGPO [41], enjoying a high-level of generalization. Specifically, the algorithm is featured by:
  • Maintenance of a replay buffer and a baseline buffer that respectively stores the trajectories for good experience replay and estimates the current policy’s average performance;
  • Recalibration of the previous advantages for off-policy update;
  • Regularization against the pre-mature entropy collapsing;
  • Shaping of the composite intrisic rewards for dominance of the outcome reward.
Compared with the vanilla GRPO-like training, the proposed method only introduced:
  • Additional policy update iterations positively associated with the number of $N_{\mathcal{D}}$ in terms of computational complexity;
  • A replay buffer of the size $N_{\mathcal{D}}$ and a baseline performance buffer of the size $N_{\mathcal{D}_{R}}$ in terms of space complexity.

### v3

@@ 4.3 Intrinsic Reward Shaping
To provide a strong baseline, we refer to the existing studies [76, 77, 78, 18] for diverse exploration, stable convergence, and effective training. Our baseline, Dr.BoT, consists of bag-of-tricks modifications to the GRPO (see Appendix A.8).
Table 1: Results on ALFWorld & WebShop (%). PT & FW stand for prompting & framework.
@@
5 Experiments
5.1 Experimental Setup
Three benchmarks are used: ALFWorld [5], WebShop [41], and DAPO-MATH-17K [46] (Appendix A.10). According to these benchmarks, we respectively follow [40] and [35] to report a range of competitive baselines (Appendix A.11). All the training settings and hyper-parameters are detailed in Appendix A.12.
5.2 Performance
Table 1 demonstrates our effectiveness on ALFWorld and WebShop. It is compatible with GRPO [45], GiGPO [40], and our Dr.BoT. SPEAR brings consistent gains across 1.5B and 7B models up to 20%. Such generalization benefits from the collection of successful trajectories, which acts as a walkthrough guide to the agent. Especially for tasks where the success rate is fairly low at the beginning, the agent has to figure out the underlying interaction logics and summarize action plans tailored specific to each task. The experience replay expedites the accumulation of tactics and thereafter reduces blind trials and errors. Furthermore, our Dr.BoT boosts GRPO and GiGPO up to 15%, showcasing the validity of mixture of tricks.
Table 2: Results (mean@30) on AIME 2024 & 2025 (%). : Official implementation already utilizes DAPO tricks. : Official results reported by Qwen [79]. PT stands for prompting.
@@ 5.2 Performance
Table 2 reports the performance of CI-integrated reasoning on AIME24 and AIME25. Dr.BoT indeed outperforms recent RL baselines. The reduced context length of Qwen3 impedes complete reasoning and answer parsing. The agent learns to exploit the CI feedback for double-check and self-reflection. SPEAR achieves comparable performance with Qwen3 but using a much smaller token budget. When the context is relaxed to 32K, an improvement is observed on both Qwen2.5 and Qwen3, confirming our generalization with more interactions turns and reasoning tokens.
@@ 5.3 Ablation Study
Table 3: Ablation on ALFWorld & WebShop. SI & IR stand for Self-Imitation & Intrinsic Reward.
@@
Table 4: Ablation on AIME 2024 & 2025 (%). SI & IR stand for Self-Imitation & Intrinsic Reward.
@@
Self-Imitation.
@@
The SIL improves baselines consistently across model scales (Table 3). Since either 1.5B or 7B models perform poorly at the early stage (i.e., success rate $15\%$), past experiences are quite beneficial to explore promising strategies. The re-use of trajectories facilitates convergence and prevents mechanical trials especially for small agents. Table 4 shows that AIME24 dropped a bit by self-imitation but AIME25 still gets improved. Such fluctuation is related to the phenomenon (Figure 4) where the imitation of samples with multiple tool calls leads to rapid increase of interaction turns and thereafter causes training instability. The competition between different reward terms affects the robust selection of good experience, ultimately degrading the effectiveness of SIL.
Intrinsic Reward.
@@
The rewarding of interaction turns benefit 1.5B models consistently (Table 3). Two 7B outliers are found where the self-imitation alone brings the most performance gains. Such exception might be related to both the task definition and the RL algorithm. One should experiment with different combinations in practice. Table 4 shows that the intrinsic reward is indispensable for both Qwen2.5 and 3 because it encourages transformation from text-based reasoning to tool-integrated reasoning. It promotes frequent tool calling and such rich observation signals motivate the agent to correct coding errors, check the validity of the answer, and reflect on alternative solutions.
@@ 5.4 Generalization on Vision-Language Agents
Figure 5: The agent learns to push the box.
To test whether the proposed SPEAR is still complimentary to existing GRPO-like algorithms on training visual agents, we follow [40] to conduct experiments on the popular visual game Sokoban [80]. In this setting, the Qwen2.5-VL-3B-Instruct [81] is adopted as the agentic LLM to solve the puzzle game where the player must push the boxes along the grid towards target positions without hitting the walls. It challenges the agent on spatial comprehension and long-term planning capabilities. The grid size is of $6\times 6$ and the visual agent receives both the visual (RGB arrays) and textual inputs as states. As shown in Table 5, the proposed method generally improves the performance on Sokoban with either GRPO, GiGPO, and the proposed Dr.BoT baselines. At first, the visual agent is unaware of the winning logic behind the game and wanders around for ”aimlessly” exploration (see Figure 5). After optimization, it not only comprehends the spatial relationship to control the box but also learns to stop moving when the task is completed.
5.5 Generalization on Search-augmented QA Tasks
Table 6: Results on search-augmented QA Tasks.
Type Method Single-Hop QA Multi-Hop QA Avg.
NQ TriviaQA PopQA HotpotQA 2Wiki MuSiQue Bamboogle
Qwen2.5-7B-Instruct
RL Search-R1 39.3 61.0 39.7 37.0 40.1 14.6 36.8 38.5
RL SPEAR 35.7 62.7 34.5 46.9 43.4 17.2 44.8 40.7
Qwen2.5-14B-Instruct
RL Search-R1 48.8 67.7 48.2 45.5 47.0 21.1 51.6 49.1
RL SPEAR 47.6 69.3 47.8 48.5 48.8 26.7 56.6 49.3
To evaluate the performance of SPEAR on knowledge-intensive reasoning tasks, we conduct experiments on search-augmented QA tasks, including the single-hop QA datasets (NQ [82], TriviaQA [83], and PopQA [84]) and multi-hop QA datasets (HotpotQA [85], 2Wiki [86], MuSiQue [87], and Bamboogle [88]). We follow the experimental settings of SearchR1 [11, 49] to launch the local wiki-18 retrieval service. We adopt the Hierarchical Navigable Small World (HNSW) CPU indexing as approximation of nearest neighbor retrieval. Our SPEAR with GRPO improves over the Search-R1 baseline on average, especially on the multi-hop QA benchmarks. Such multi-hop QA datasets require reasoning for problem decomposition and several turns of information seeking. In this case, our intrinsic reward that encourages multiple tool uses for broad exploration prevents arbitrary conclusions with only one or two searches. Our SPEAR respectively requires $\sim 14.48$ and $\sim 14.42$ calls for 7B and 14B models, respectively. Such behavior is expected due to the stimulation of exploration at the beginning. Despite the QA tasks are relatively short-horizon, the agent still benefits from the detailed decomposition of the complex queries with cross-validation via step-by-step searching. Note that our retrieval service adopts the HNSW of E5 embedding for efficient training, which slightly impedes performance [89].
5.6 More Discussions
Due to the page limit, discussions on theoretical analysis on convergence A.9, hyper-parameters A.13, qualitative analysis A.14, training cost and complexity A.15, and future research directions A.16 are presented in the appendix. One could easily adapt SPEAR to training any (M)LLM-driven agents robustly without binding to a specific optimization algorithm.
@@ 6 Conclusions and Limitations
In this paper, we target the pivotal challenge of balancing exploration and exploitation in RL training of LLM agents. Our proposed solution, SPEAR [Uncaptioned image], extends the vanilla SIL by advantage recalibration, scheduled entropy control, and intrinsic rewards. These components work in a curriculum manner to prevent policy collapse and excessive uncertainty, progressively guiding the policy through a smooth transition between exploration and exploitation. In addition, we propose a strong baseline Dr.BoT tailored for agentic RL with existing bag-of-tricks verified from numerical industrial practices. Empirical results across tasks and models showcase SPEAR’s superiority over existing methods, with performance boosts and acceptable computational overhead. The effectiveness of our SPEAR underscores the value of learning from past experiences while managing policy entropy, offering a robust framework for training LLMs with strong reasoning and tool integration skills.
There exist two potential limitations: 1) The vague definition of good experiences under highly complex, stochastic environments with unreliable tools. In such cases, observations can be noisy and severely degrade the feasibility of the task. The sparse outcome reward cannot distinguish between good and bad experiences and thereafter the relative advantages might be simply attributed to randomness instead of the agent’s behavior. We suggest a possible solution that more fine-grained, stepwise supervision should be enforced. For example, a step-wise process reward that evaluates the logical consistency [90] of the agent’s thought and action might be helpful. 2) The rigidity of entropy control which relies on prior-based scheduling and covariance-based clipping. In the present study, the proposed scheduling and clipping designs might not be optimal for all kinds of agentic tasks. A more adaptive solution lies in the policy’s self-confidence on decisions under each observation. One might use the token-level dynamic reweighting for SIL [91] which avoids over-concentration on certain low-probability reference tokens in the replay buffer. Similarly, the clipping could depend on token probability instead of the bounded random sampling. We leave the exploration mentioned above as a promising direction for improvement in the future.
@@ Appendix A Appendix
A.1 Summary of the Appendix
In the appendix, we provide detailed explanations on the following.
  • Descriptions about the Action Space
  • Brief Introduction to the PPO and GRPO
  • PseudoCode of the SPEAR
  • Visualization of the Curriculum Schedule
  • Definition of the Reward Function
  • Descriptions about RL Bag-of-Tricks
  • Theoretical Analysis on Convergence
  • Descriptions of the Data and Environment
  • Choice of Baselines
  • Implementation Details
  • Discussions and Guidelines on Hyper-parameters
  • Qualitative Analysis
  • Training Cost and Complexity
  • Future Research Directions
A.2 Detailed Action Space
The following contents correspond to Section 3.1 in the main text.
TextWorld Embodied Tool.
The embodied actions follows ALFWorld [5] where a language-driven agent interacts with the TextWorld [92]. It allows the agent to take one of the following high-level actions: goto {recep}, take {obj} from {recep}, put {obj} in/on {recep}, open {recep}, close {recep}, toggle {obj}{recep}, clean {obj} with {recep}, heat {obj} with {recep}, and cool {obj} with {recep}, where {obj} and {recep} denote objects and receptacles, respectively.
Web Browsing Tool.
The definition of web browsing follows WebShop [41] where only two actions are allowed: search[query] and choose[button] where query and button respectively stand for searching query and clickable elements such as back to search, prev/next page, {product title}, {option}, {desc/overview}, previous, and buy.
Code Interpreter Tool.
The code interpreter executes the code generated by the language model and return both the stdout and stderr. If the code runs correctly, the stdout contains the output. On the other hand, the compiler error messages are provided for the next-round correction. We follow [35] to deploy a SandBox [93] service that receives execution requests from the interpreter tool. In addition, we add a reminder in the stdout for empty output when the LLM forgets to print computation results: Empty stdout! You might forget to print the answer. For non-empty stderr, we also add an instruction as hint: Errors occurred! Check your code.
A.3 Detailed Policy Optimization Algorithms
The following contents correspond to Section 3.2 in the main text.
Proximal Policy Optimization (PPO).
Typically, PPO optimizes the following:
Group Relative Policy Optimization (GRPO).
A.4 Pseudo Code
The following contents correspond to Section 4 in the main text.
Algorithm 1 summarizes the full training procedure of the proposed SPEAR. It is noted that our SPEAR is compatible with various baselines such as GRPO [45] and GiGPO [40], enjoying a high-level of generalization. Specifically, the algorithm is featured by: 1) Maintenance of a replay buffer and a baseline buffer that respectively stores the trajectories for good experience replay and estimates the current policy’s average performance; 2) Recalibration of the previous advantages for off-policy update; 3) Regularization against the pre-mature entropy collapsing; 4) Shaping of the composite intrisic rewards for dominance of the outcome reward.
Compared with the vanilla GRPO-like training, the proposed method only introduced: 1) Additional policy update iterations positively associated with the number of $N_{\mathcal{D}}$ in terms of computational complexity; 2) A replay buffer of the size $N_{\mathcal{D}}$ and a baseline performance buffer of the size $N_{\mathcal{D}_{R}}$ in terms of space complexity.